A worker , tearing down a barn , uses a pinch bar 4ft long with the shorter end 2 in. from the fulcrum. If the worker can exert a force of 150lbs . at one end of the pinch bar, how great a force can he exert at the other end?(Note the differences in units, 2 in and 4 ft . Units must be the same when using the lever formula.)
Answer by checkley75(3666)

2/150=48/X CROSS MULTIPLY
2X=48*150
2X=7200
X=7200/2
X=3600 LBS.